# miniCTX: Neural Theorem Proving with

### v1

@@ Abstract
We introduce miniCTX, which tests a model’s ability to prove formal mathematical theorems that depend on new definitions, lemmas, or other contextual information that was not observed during training. miniCTX contains theorems sourced from real Lean projects and textbooks, each associated with a context that can span tens of thousands of tokens. Models are tasked with proving a theorem given access to code from the theorem’s repository, which contains context that is helpful or needed for the proof. As a baseline for miniCTX, we introduce file-tuning, a simple recipe that trains a model to generate a proof step conditioned on the preceding file contents. File-tuning substantially outperforms the traditional neural theorem proving approach that fine-tunes on states alone. Additionally, our file-tuned model improves performance on the standard miniF2F benchmark, achieving a pass rate of 33.61%, which is a new state-of-the-art for 1.3B parameter models. Alongside miniCTX, we offer ntp-toolkit for automatically extracting and annotating theorem proving data, making it easy to add new projects into miniCTX to ensure that contexts are not seen during training. miniCTX offers a challenging and realistic perspective on evaluating neural theorem provers. [^1] Data and models: HuggingFace. ntp-toolkit: cmu-l3/ntp-toolkit. Evaluation code: cmu-l3/minictx-eval.
@@ 1 Introduction
Formal theorem proving in interactive theorem provers (ITPs) provides a testbed for evaluating the reasoning capabilities of large language models (LLMs). Theorem proving capabilities can then directly translate to automation for mathematicians, such as via tools that complete or formalize proofs [1, 2, 3, 4]. However, despite their promise, we see a gap between current language model-based provers and the complexity of real-world theorem proving.
@@
Our motivating observation is that theorems and proofs depend on various forms of context, such as newly-defined definitions and lemmas. For instance, to prove results about a square, one might first formalize a definition of a rectangle, prove some results about rectangles, then specialize them to a newly-defined square [5] (Figure 1). However, existing methods for training and evaluating LLM-based theorem provers do not take such context into account. For example, benchmarks often focus on proving standalone competition problems (e.g., miniF2F [6]) or theorems from a library that the model has trained on (e.g., mathlib [7, 8]), and LLM-based provers are trained to accept only a proof state as input, making them unaware of new theorems and definitions [9].
From a language modeling perspective, a key challenge is the sheer size of potentially useful context. For instance, proving a theorem in the Prime Number Theorem project without having seen it during training can require reasoning over a long context of interdependent definitions and theorems [10]. Therefore, theorem proving with newly-defined contexts offers a challenging testbed for long-context reasoning, in addition to more accurately reflecting practical proving. With these considerations in mind, we define a new setting of context-dependent theorem proving, and develop an evaluation framework, benchmark, and training recipe for theorem proving with context.
@@ 2 Theorem proving with context
Formal theorem proving involves two stages: defining mathematical objects and facts relevant to the desired result, then stating and proving the result itself. For language model-based provers to function as useful tools in this real-world setting, they need to be able to work with new information such as new definitions or lemmas. For example, a system suggesting proofs in the Prime Number Theorem project should be familiar with the project’s definition of “square”. Unfortunately, current language model-based provers are trained on a static snapshot of data, and are hence unaware of any context that was created after the model was trained (Figure 1). As a result, it remains unclear whether these models can work with new information, and how to enable this capability.
Figure 1: Traditional language model-based provers train on a static collection of data (e.g., a snapshot of Mathlib), and are evaluated on standalone theorems that depend only on facts and theorems seen during training (e.g., competition problems [6] that rely only on basic facts in algebra and number theory). We argue that neural theorem provers must also operate in the realistic context-dependent setting, in which results depend on working with new mathematical objects and their facts (right).
Context-dependent proving.
We study context-dependent theorem proving, where the goal is for a model to generate proofs $y$ for new theorems $x$, based on a context $c$ that includes background information such as definitions, lemmas, or natural language comments. Formally,
$\arg\max_{M}\mathbb{E}_{(x,c)\sim p}\mathbb{E}_{y\sim M(\cdot|x,c)}v(x,c,y),$ (1)
where $(x,c)\sim p$ denotes a (theorem, context) pair from a context distribution $p$, $M$ is a model that produces a proof $y$, and $v$ returns 1 if the proof is correct and 0 otherwise (implemented by Lean [^2] We focus on the Lean [11] proof assistant in this paper, though the problem of context-dependent proving is present in other proof assistants such as Coq [12] and Isabelle [13].).
In this paper, we treat a Lean repository as a context distribution. Each context $c$ is some subset of code in the repository that does not contain the theorem’s formal proof. In this paper we take $c$ to be the code that precedes the theorem $x$ in a file. This context includes new definitions, lemmas, and comments that are relevant to the theorem but does not show the theorem’s formal proof.
Evaluating context-dependent proving.
To measure a model’s performance on context-dependent proving, a model is tasked with generating a proof $y$ given a theorem $x$ and its preceding code $c$. Given a language model, we can test three kinds of generalization by ensuring the following:
@@
Assuming that we know the date at which a model was trained, we can guarantee theorem-level, context-level, and project-level generalization by selecting theorems, files, or projects (respectively) that were added after the model’s training date.
To realize our evaluation framework, we develop a method that automatically extracts a dataset of (context, theorem, metadata) examples from a Lean repository, i.e., $f(R)\rightarrow\{(c,x,z)_{i}\}_{i=1}^{N}$. Next, we apply this method to create miniCTX, a benchmark for context-dependent theorem proving.
@@
We develop miniCTX, a Lean 4 theorem proving benchmark of theorems that depend on newly-defined lemmas, definitions, and proofs from within a project. miniCTX is currently based on 376 theorems from four projects: recent Lean formalizations of the Prime Number Theorem (PrimeNumberTheoremAnd[10] and the Polynomial Freman-Ruzsa Conjecture (PFR[14], problems from an introductory text on theorem proving (HTPI[15], and recent results from Lean’s mathematical library (Mathlib[16].
Each theorem in miniCTX is accompanied by the following data, formatted in JSON:
@@
      Commit at which the theorem and its file was added,
@@
      Position of the theorem and number of premises preceding it,
@@
      Proof length and type,
@@
      Whether the statement or proof uses definitions or lemmas from within the file or repository.
See example in §A.1. We make the dataset available at l3lab/miniCTX, along with corresponding evaluation code based on the Lean REPL [17]. The dataset has an Apache 2.0 license.
Prime Number Theorem.
PrimeNumberTheoremAnd [10] is a project started in January 2024 that formalizes the prime number theorem in Lean as well as related concepts, such as residue calculus on rectangles in $\mathbb{C}$. We find the files Rectangle.lean and ResidueCalcOnRectangles.lean suitable for our purpose of testing context-dependent theorem proving, especially when we use preceding file content as context, as each file is self-contained within the project and contains new definitions (rectangles, squares) and many interdependent lemmas. See §A.2 for an illustration of such lemmas. In addition, most theorems from ResidueCalcOnRectangles.lean rely on the definitions from Rectangle.lean, which serves as a perfect example of cross-file dependencies. We extracted 87 theorems from these files. Assuming that a model was trained prior to January 2024, this split guarantees the evaluation of project-level, context-level, and theorem-level generalization.
PFR.
PFR [14] is a project started in November 2023 that formalizes a proof of the Polynomial Freiman–Ruzsa (PFR) conjecture. We included 54 theorems from PFR. We find that proofs of theorems in PFR tend to be much more monolithic and longer in length than those in Mathlib or other libraries. PFR also defines custom mathematical concepts and notations (such as Ruzsa distance) and a proof typically depends on many lemmas in PFR outside the current file. All of the theorems were added to PFR after November 2023. Assuming that the model was trained prior to this date, this split guarantees the evaluation of project-level, context-level, and theorem-level generalization.
Recent Mathlib Commits.
Lean’s mathematical library, Mathlib [16], is a community-maintained Lean repository including mathematical concepts as well as programming APIs and common tactics. It is the single largest Lean library that users contribute to, and is therefore representative of the production environment in which neural theorem provers are deployed. Mathlib is a long-standing project, and it is common practice to train language model-based provers on Mathlib. It is therefore likely that Mathlib source files have been observed during training. However, Mathlib is frequently updated, with new definitions, theorems, and refactorings occurring on a daily basis. Hence we can test theorem-level generalization by tasking a model with proving newly added theorems, given a context that may or may not have been observed during training.
We included 50 theorems added to Mathlib in April 2024, by filtering recent Mathlib commits to ones that only add new theorems. Many of the theorems added are simple lemmas that depend on earlier ones (e.g., ones seen during training). As Mathlib generally refactors new theorems by breaking down long theorems to shorter lemmas, most new theorems are not difficult to prove, and give a realistic representation of where neural theorem provers are used. Assuming that the model was trained prior to April 2024, the Mathlib split guarantees the evaluation of theorem-level generalization.
HTPI.
HTPI contains the Lean code for the book How to Prove It (HTPI) [18], which explains a systematic approach to constructing mathematical proofs with Lean. It covers various topics, including elementary logic, number theory, and proving techniques like mathematical induction, along with their implementation in Lean. As supplementary material to the textbook, the problems in HTPI are formulated in a similar fashion: the files typically start with basic definitions and lemmas that might be used throughout the entire file, followed by exercises and several example problems. [^3] The main chapter and exercises are separated in the original project: HTPILeanPackage. We manually merged them for evaluation in the fork: HTPILeanPackage4.7. Therefore, models can utilize definitions, lemmas, and proof structures from example problems to solve similar exercises. Intuitively, the model must understand and apply the context provided within each file, making it an effective benchmark for testing context-aware theorem-proving models.
Split Problems Avg. Context Length (tokens) Avg. Proof Steps
miniF2F [[6]] Valid/Test 488     153*    3.0**
miniCTX Prime 87   10,630    3.6
PFR 54   17,495 27.7
@@
HTPI 185   39,050    10.7**
All 376   26,106 10.9
* Average context length for MiniF2F only counts libraries and definitions.
** Average proof steps for MiniF2F and HTPI only include problems with proofs.
Table 1: Overview of Problem Statistics in miniF2F and miniCTX.
3.1 Core challenges: dependencies, length, constraints, and varied content
miniCTX introduces several challenges for language-model based provers.
File and repository dependencies. First, a theorem’s statement or proof may have contextual dependencies, i.e., requiring the use of definitions or lemmas from the new repository. Assuming that the model has not observed the repository during training, a model must have some mechanism for taking in the new definitions and lemmas, and must have the ability to produce proofs that, intuitively, require understanding the new context. As a proxy indicator of having dependencies, we record whether the statement or proof uses definitions or theorems from within the file or repository.
Long context length.
A second challenge is the length of the context in miniCTX. As seen in Table 1, the average context length is over 26,000 tokens. Intuitively, the model must reason over long contexts containing information that is needed to produce a proof. This differs from proving standalone competition problems in miniF2F, which have an average context length of 153 tokens.
Constraints. Third, real-world proofs may have particular constraints, such as disabling common automation tools and defining new tactics, or custom definitions and syntax such as in PFR. A model must be able to adapt and produce these context-dependent proofs.
Varied content and difficulty. A final challenge is that miniCTX has varied mathematical content in its theorems, thus requiring a flexible automated proving system. This ranges from writing explicit calculation proofs in HTPI to writing sophisticated proofs in PFR. One axis of variation is the length of the human-written proof (if available), which we record and treat as a proxy for difficulty.

### v2

@@ Abstract
Real-world formal theorem proving often depends on a wealth of context, including definitions, lemmas, comments, file structure, and other information. We introduce miniCTX, which tests a model’s ability to prove formal mathematical theorems that depend on new context that is not seen during training. miniCTX contains theorems sourced from real Lean projects and textbooks, each associated with a context that can span tens of thousands of tokens. Models are tasked with proving a theorem given access to code from the theorem’s repository, which contains context that is needed for the proof. As a baseline for miniCTX, we tested fine-tuning and prompting methods that condition theorem proving on preceding context. Both approaches substantially outperform traditional methods that rely solely on state information. We found that this ability to use context is not captured by previous benchmarks such as miniF2F. Alongside miniCTX, we offer ntp-toolkit for automatically extracting and annotating theorem proving data, making it easy to add new projects into miniCTX to ensure that contexts are not seen during training. miniCTX offers a challenging and realistic evaluation of neural theorem provers. [^1] Data and models: HuggingFace. ntp-toolkit: cmu-l3/ntp-toolkit. Evaluation code: cmu-l3/minictx-eval.
@@ 1 Introduction
Formal theorem proving in interactive theorem provers (ITPs) provides a testbed for evaluating the reasoning capabilities of large language models (LLMs). Theorem proving capabilities can then directly translate to automation for mathematicians, such as via tools that complete or formalize proofs (Welleck & Saha, 2023; Song et al., 2024; Welleck, 2024; Agrawal et al., 2022). However, despite their promise, we see a gap between the evaluation of current language model-based provers and the complexity of real-world theorem proving.
@@
Our motivating observation is that theorems and proofs depend on various forms of context, such as newly-defined definitions and lemmas. For instance, to prove results about a square, one might first formalize a definition of a rectangle, prove some results about rectangles, then specialize them to a newly-defined square (Kontorovich, 2024b) (Figure 1). However, existing methods for training and evaluating LLM-based theorem provers often fail to incorporate the full range of contextual information available in real-world projects. For example, benchmarks often focus on proving standalone competition problems (e.g., miniF2F (Zheng et al., 2022)) or theorems from a library that the model has trained on (e.g., Mathlib (Han et al., 2022; Yang et al., 2023)), and state-of-the-art LLM-based provers are trained to accept only a proof state as input, making them unaware of new theorems and definitions (Polu & Sutskever, 2020; Ying et al., 2024; Xin et al., 2024). While some existing work, including premise selection techniques (Jiang et al., 2022; Mikuła et al., 2023; Yang et al., 2023) and datasets like CoqGym (Yang & Deng, 2019), have explored theorem proving based on information beyond the current state, they often focus only on providing relevant premises—lemmas that can assist proof construction—which are only a subset of the available information.
Building on these foundations, we propose miniCTX: a benchmark that seeks to expand the scope of context used in theorem proving. We extend beyond traditional premise selection explored in prior benchmarks (e.g., Yang et al. (2023); Yang & Deng (2019)) by incorporating a more comprehensive set of contextual elements. This includes premises, prior proofs, comments, notation, and structural components like imports and declarations. By doing so, miniCTX aims to drive the development of methods that understand and work with context that occurs in complex, real-world theorem proving tasks. We compare miniCTX with several popular theorem proving datasets to highlight its unique contributions in terms of contextual dependency and real-world applicability (see Table 1). Additionally, considering the common use of pre-trained language models we mitigate potential data contamination by continually and automatically updating miniCTX with new Lean projects, so that evaluated theorems are not seen during training. Our key contributions are:
miniCTX Benchmark: We introduce miniCTX, the first benchmark designed specifically to evaluate theorem proving in real-world settings where proofs depend on in-file definitions, lemmas, and context from formal projects. miniCTX presents a unique challenge by requiring models to reason over long contexts and handle dependencies that arise in real-world theorem proving tasks.
ntp-toolkit: To facilitate the automatic updating of miniCTX, we developed the ntp-toolkit, which automatically extracts relevant theorems and contexts from Lean projects. Additionally, we provide a Lean REPL wrapper that enables simpler evaluation on miniCTX.
Baseline Evaluations: We evaluate miniCTX on several existing baseline models, including different fine-tuning and prompting strategies, as well as methods with premise selection. We also propose file-tuning, a strong baseline method for training models using full file contexts, where both the theorem statements and their surrounding context are provided during training. This approach establishes a robust baseline for future work on context-dependent theorem proving.
Table 1: Comparison of theorem proving benchmarks across several key features.
Benchmark Language Premise Full Context Multi-source Temporal Split
miniF2F (Zheng et al., 2022) Multiple
ProofNet (Azerbayev et al., 2023) Lean
LeanDojo (Yang et al., 2023) Lean
LeanStep (Han et al., 2022) Lean
CoqGym (Yang & Deng, 2019) Coq
PISA (Jiang et al., 2021) Isabelle
miniCTX (Ours) Lean
Figure 1: Many state of the art provers are trained on a static dataset of theorems and proofs, then evaluated on standalone problems such as competition problems (left). We argue that neural provers must also operate in the realistic context-dependent setting, in which results depend on working with new mathematical objects and their facts, notations, and the structural elements of the project (imports, variables, etc.) (right).
@@ 2 Theorem proving with context
For language model-based provers to function as useful tools in this real-world setting, they need to be able to work with new information such as new definitions or lemmas. For example, a system suggesting proofs in the Prime Number Theorem project should be familiar with the project’s definition of “square”. Many current language model-based provers are trained on a static snapshot of data, and are hence unaware of any context that was created after the model was trained (Figure 1). They are often evaluated on either context-less standalone competition problems (Zheng et al., 2022; Azerbayev et al., 2023) that do not reflect realistic settings, or existing Mathlib proofs (Yang et al., 2023) which risks data contamination. As a result, it remains unclear whether these models can work with new information, which is necessary for using them as an assistant in a real proof development, and how to enable this capability.
Context-dependent proving. We study context-dependent theorem proving, where the goal is for a model to generate proofs $y$ for new theorems $x$, based on a context $c$ that includes background information such as definitions, lemmas, or natural language comments. Formally, the problem is
$\mathrm{maximize}_{M}\,\mathbb{E}_{(x,c)\sim p}\mathbb{E}_{y\sim M(\cdot|x,c)}v(x,c,y),$ (1)
where $(x,c)\sim p$ denotes a (theorem, context) pair from a context distribution $p$, $M$ is a model that produces a proof $y$, and $v$ returns 1 if the proof is correct and 0 otherwise.
Evaluating context-dependent proving. We choose Lean (Moura & Ullrich, 2021) as the verifier $v$, because of the large body of recent theorems in Lean that can be used as evaluation data, and the abundance of proving methods in Lean that we use as baselines. We treat a Lean repository as the distribution $p$. Each context $c$ is a subset of the repository, including new definitions, lemmas, notations, imports, and comments that are relevant to the theorem.
Given a language model, we can test three kinds of generalization by ensuring the following:
@@
For baseline evaluations, we investigate the in-file context case where $c$ is the source code that precedes the theorem $x$ in a file, as well as cross-file context where $c$ includes both preceding code and relevant premises retrieved from imported modules through premise selection.
@@
We develop miniCTX, a Lean 4 theorem proving benchmark of theorems that depend on newly-defined lemmas, definitions, and proofs from within a project. miniCTX is currently based on 376 theorems from six projects: (1) Prime Number Theorem (PNT) (Kontorovich, 2024a), (2) Polynomial Freiman-Ruzsa Conjecture and its extension with high cross-file dependency (PFR) (Tao, 2023), (3) recent results from the standard mathematical library (Mathlib), (4) an introductory text on theorem proving (HTPI) (Macbeth, 2023), (5) high energy physics formalization in HepLean (HEP) (Tooby-Smith, 2024), and (6) scientific computing formalizations (SciLean) (Skřivan, 2021). Table 2 shows an overview of the dataset. Each theorem in miniCTX consists of the theorem statement, preceding file contents up to the theorem statement, and metadata, in JSON (see §A.1).
@@
      Commit and time at which the theorem and its file was added,
@@
      Position of the theorem in the file and number of premises preceding it,
@@
      Number of in-file premises and cross-file premises used by the statement or proof,
@@
      Imported modules, along with separate files recording the premises defined in each imported module,
    7. (g)
      Proof length and type.
Using our benchmark, users can easily reconstruct the complete context for each theorem, including both in-file and cross-file context. The in-file context is provided directly by preceding file contents, while the cross-file context can be reconstructed using the metadata, which includes information on imported modules. We open-source the dataset and evaluation code.
Table 2: Problem statistics in miniF2F (Zheng et al., 2022) and miniCTX.
Split Problems Context Size In-File Premises Repo. Premises Proof Size
(tokens) (premises / 100 tokens) (premises / 100 tokens) (lines)
miniF2F Valid/Test 488   153*  3.0
miniCTX PNT 85   10,858 1.87 0.30   3.3
PFR 54   18,059 0.65 1.10 27.2
PFRcross 43   4,351 0.44 2.75   2.7
@@
HTPI 45   65,082 2.85 0.00 10.7
Hep 61   3,585 5.65 4.25   3.1
SciLean 46   6,249 2.08 9.72   1.8
All 384   18,690 1.94 2.63    8.5
*Only counting library imports and definitions.   Excluding theorems without proofs.
3.1 miniCTX Sources
PNT. PrimeNumberTheoremAnd (Kontorovich, 2024a) is a project started in January 2024 that formalizes the prime number theorem in Lean as well as related concepts, such as counter integral on rectangles in $\mathbb{C}$. We find the files Rectangle.lean and ResidueCalcOnRectangles.lean suitable for our purpose of testing context-dependent theorem proving, especially when we use preceding file content as context, as each file is self-contained within the project and contains new definitions (rectangles, squares) and many interdependent lemmas. See §A.2 for an illustration of such lemmas.
PFR. PFR (Tao, 2023) is a project started in November 2023 that formalizes a proof of the Polynomial Freiman–Ruzsa (PFR) conjecture. We included 54 theorems from PFR. We find that proofs of theorems in PFR tend to be much more monolithic and longer in length than those in Mathlib or other libraries. PFR also defines custom mathematical concepts and notations (such as Ruzsa distance) and a proof typically depends on many lemmas in PFR outside the current file.
PFR_crossfile. PFR_crossfile is an extension of the PFR split, which includes additional problems to further evaluate cross-file dependencies. To evaluate models’ performance on problems with extensive cross-file dependencies, we added 43 theorems from the files Entropy.Group and Entropy.Kernel.Group, which have the highest cross-file dependencies in the project. These problems contain three times the number of cross-file premises compared to other math splits, making them a strong candidate for evaluating a model’s ability to utilize cross-file premises effectively.
Recent Mathlib Commits. Mathlib (Mathlib Community, 2020), is Lean’s largest community-maintained repository, encompassing a wide range of mathematical concepts, programming APIs, and common tactics. It is commonly used for training theorem-proving models and is frequently updated with new definitions, theorems, and refactorings. Therefore, to avoid data contamination, we included 50 theorems newly added to Mathlib in April 2024, by filtering recent Mathlib commits to ones that only add new theorems. Assuming that the model was trained prior to April 2024, the Mathlib split guarantees the evaluation of theorem-level generalization.
HTPI. HTPI contains the Lean code for the book How to Prove It (HTPI) (Velleman, 2019), which explains a systematic approach to constructing mathematical proofs with Lean. It covers topics like elementary logic and number theory, and proving techniques like induction.The files in HTPI typically start with basic definitions and lemmas that might be used throughout the entire file, followed by exercises and several example problems. Therefore, models can utilize definitions, lemmas, and proof structures from example problems to solve exercises, making it an effective benchmark for testing context-dependent theorem-proving models.
HEP. HepLean (Tooby-Smith, 2024) is an open-source project that digitalizes definitions, theorems, proofs, and calculations in high energy physics using Lean. HepLean aims to facilitate discovery, automate new findings, verify correctness, and provide educational tools in physics. We selected files from the space and time section, which introduces several new definitions, including 4D Euclidean spacetime models. These files contain high in-file and cross-file dependencies, ideal for evaluating context-dependent theorem proving. This split represents an area outside of mathematics, further expanding the generalization of the benchmark.
Recent SciLean Commits. SciLean (Skřivan, 2021) is an open-source project designed to formalize concepts in scientific computing using Lean 4. SciLean aims to provide formalized tools and frameworks for efficiently representing and proving properties of numerical methods, differential equations, and optimization problems, bridging applied mathematics and computer science. Given Scilean has been created for three years and accessed by multiple projects, similar to Mathlib split, we selected 46 theorems added to Scilean in April 2024 in order to ensure the testing data is more likely unseen for evaluating models. The newly added problems are mostly supplements to existing files, providing an opportunity to evaluate the model’s ability to apply and learn from previous lemmas.
3.2 Problem selection methodology
The problem selection process for different splits in miniCTX follows three main criteria: (1) ensuring new problems are less likely to have appeared in training data, (2) utilizing an automated selection process, and (3) promoting generalization. Depending on the properties and goals of each split, we adopted three primary approaches:
  1. 1.
    Recency-based selection: For popular libraries such as Mathlib and SciLean, which are highly likely to be used for training, we aim to select newly added theorems. This is achieved by sorting theorems based on the timestamp of when the theorem was first added to the project, which is extracted by ntp-toolkit. This strategy helps mitigate data contamination risks.
  2. 2.
    Random selection: For more recent projects, such as the PFR split, where the risk of data contamination is lower, we randomly select proved theorems to ensure a representative sample of the entire project. This approach maintains generality of the selected problems.
  3. 3.
    Dependency-based selection: To explicitly evaluate models’ performance in context-dependent proving, we selected files based on the in-file and cross-file premise labels available in our benchmark. For the Prime split, we chose the file with the highest number of in-file dependencies, while for PFRcrossfile, we selected files with the highest cross-file dependencies. For HEPLean, we selected files with a balance of both types. The level of dependency is also extracted automatically by ntp-toolkit.
Although human inspection and expertise are involved in ensuring that the selected problems are valid and sufficiently general for evaluating models across diverse settings, all selection processes are ultimately automated by using the labels extracted through our toolkit. This ensures consistency across the benchmark and scalability to future updates.
3.3 Key features and challenges
miniCTX introduces several key features that distinguish it from other theorem proving benchmarks, addressing challenges that have not been tackled by previous benchmarks:
Real-world theorem proving. Unlike popular benchmarks (e.g., miniF2F (Zheng et al., 2022), ProofNet (Azerbayev et al., 2023), FIMO (Liu et al., 2023)) that focus on isolated competition problems, real-world research-level theorem proving is heavily dependent on rich mathematical contexts. Therefore, miniCTX includes real-world, complex theorems from a variety of ongoing Lean projects, such as Prime Number Theorem (Prime) and Polynomial Freiman-Ruzsa Conjecture (PFR). They rigorously test a model’s ability in real-world formalization projects. This diversity contrasts with the LeanDojo benchmark (Yang et al., 2023), which focuses solely on Mathlib, enabling miniCTX to better test a model’s generalization in different settings.
Contextual evaluation. Proving a theorem often depends on new definitions, lemmas, or other contextual information, which a model may not have seen during training. miniCTX includes theorems along with this new context. During evaluation, the model is expected to leverage the provided new context to help prove the theorem.